# Intersection numbers as mixed volumes of Newton–Okounkov bodies

### v2

@@ 4. Construction of the flag
In this section we show the existence of an admissible flag $Y_{\bullet}$ satisfying the properties (i)–(iii) in Theorem 1.2. For simpler notation we set
@@
Proceeding by induction, let us assume that we have already constructed $Y_{j}$ for some $20\leq j\leq d-$ 2. By induction we may also assume that $Y_{j}\cap X_{\mathrm{reg}}\neq\emptyset$. Since $L^{\prime}_{j+1}|_{Y_{j}}$ is ample on $Y_{j}$, there exists an $\rho_{j+1}\in\mathbb{Q}_{>0}$ such that $\rho_{j+1}L^{\prime}_{j+1}|_{Y_{j}}\cong\iota_{j}^{*}\mathcal{O}(1)$ for some embedding $\iota_{j}\colon Y_{j}\to\mathbb{P}^{N_{j}}$. Hence, we will consider $Y_{j}$ as a subvariety of $\mathbb{P}^{N_{j}}$ in the following. We use the following two versions of Bertini’s theorem: By [6, Theorem 1.1] any general hyperplane $H\subseteq\mathbb{P}^{N_{j}}$ intersects $Y_{j}$ in an irreducible Cartier divisor on $Y_{j}$. By [5, Corollary 2] we have $(H\cap Y_{j})_{\mathrm{reg}}\supseteq Y_{j,\mathrm{reg}}\cap H$ for any general hyperplane $H\subseteq\mathbb{P}^{N_{j}}$. Since every hyperplane $H\subseteq\mathbb{P}^{N_{j}}$ intersects $Y_{j}$, any general hyperplane $H\subseteq\mathbb{P}^{N_{j}}$ intersects the dense open subset $Y_{j}\cap X_{\mathrm{reg}}$ of $Y_{j}$. Hence, there is a hyperplane $H_{1}\subseteq\mathbb{P}^{N_{j}}$ such that $Y_{j+1}=H_{1}\cap Y_{j}$ is an irreducible Cartier divisor on $Y_{j}$, every regular point of $Y_{j}$ lying in $Y_{j+1}$ is also regular in $Y_{j+1}$, and $Y_{j+1}\cap X_{\mathrm{reg}}\neq\emptyset$. By construction it holds $\rho_{j+1}L^{\prime}_{j+1}|_{Y_{j}}\cong\mathcal{O}_{Y_{j}}(Y_{j+1})$. By induction we end up with an admissible flag satisfying properties (i)–(iii) in Theorem 1.2 with $r_{j}=\rho_{j}^{-1}$ after setting $Y_{d}=\{p\}$ for any point $p\in Y_{d-1}\cap X_{\mathrm{reg}}$.
@@ 5. The Newton–Okounkov bodies in this paper
Let $L_{1},\dots,L_{d}$ be ample line bundles on $X$ and $Y_{\bullet}$ an admissible flag on $X$ satisfying the corresponding properties (i)–(iii) in Theorem 1.2 with $k\geq 1$. Then it holds
@@
    As $L_{1}\equiv r_{1}\mathcal{O}_{X}(Y_{1})$, we can compute by the slice formula (2.5), Equation (2.2) for real numbers and Equations (2.3) and (2.4)
@@
    Note for the first inclusion, that we always have $\Delta_{Y_{\bullet}|Y_{1}}(L)\subseteq\Delta_{Y_{1,\bullet}}(L|_{Y_{1}})$ for any line bundle $L$ on $X$. This completes the proof of the lemma.
@@
We will prove Theorem 1.2 by induction on $k$. By Section 4 we only have to prove the second statement. If $k=0$, we proved in [17, Theorem 1.2] that the inclusion (2.3) is indeed an equality $\Delta_{Y_{\bullet}}(N_{1})+\Delta_{Y_{\bullet}}(N_{2})=\Delta_{Y_{\bullet}}(N_{1}+N_{2})$ if $N_{1},N_{2}\in C_{L}(M)$. Thus, the theorem follows from multilinearity of the intersection number and the mixed volume, see also [17, Section 4]. Hence, we may assume $k\geq 1$. In particular, $d\geq 2$. Applying the induction hypothesis to $Y_{1}$ we get
$(L_{1}\cdot\ldots\cdot L_{d})=r_{1}(L_{2}|_{Y_{1}}\cdot\ldots\cdot L_{d}|_{Y_{1}})=r_{1}(d-1)!V(\Delta_{Y_{1,\bullet}}(L_{2}),\dots,\Delta_{Y_{1,\bullet}}(L_{d})).$
Since $L_{2},\dots,L_{d}$ are ample, we get by Equation (2.4)
$V(\Delta_{Y_{1,\bullet}}(L_{2}),\dots,\Delta_{Y_{1,\bullet}}(L_{d}))=V(\Delta_{Y_{\bullet}|Y_{1}}(L_{2}),\dots,\Delta_{Y_{\bullet}|Y_{1}}(L_{d})).$
@@
where we consider the convex bodies in the mixed volume on the right hand side as convex bodies in $\mathbb{R}^{d-1}\cong\{0\}\times\mathbb{R}^{d-1}$. By Lemmas 3.1 and 5.1 we get
@@
In this section we prove Corollary 1.3. It is enough to prove that for every $\epsilon>0$, every birational morphism $\pi\colon\widetilde{X}\to X$ for a smooth projective variety $\widetilde{X}$, and all nef line bundles $\widetilde{L}_{1},\dots,\widetilde{L}_{d}$ on $\widetilde{X}$ such that $E_{i}=\pi^{*}L_{i}-\widetilde{L}_{i}$ is effective, there is a rank $d$ valuation $\nu\colon K(X)^{\times}\to\mathbb{Z}^{d}$ with
(7.1) $(\widetilde{L}_{1}\cdot\ldots\cdot\widetilde{L}_{d})\leq d!V(\Delta_{\nu}(L_{1}),\dots,\Delta_{\nu}(L_{d}))+\epsilon.$
We choose an ample line bundle $A$ on $\widetilde{X}$ and a positive rational number $\delta>0$ to be determined later. Let $\widetilde{\nu}=\nu_{Y_{\bullet}}\colon K(\widetilde{X})^{\times}\to\mathbb{Z}^{d}$ be a valuation obtained from an admissible flag $Y_{\bullet}$ on $\widetilde{X}$ associated to the ample $\mathbb{Q}$-line bundles $\widetilde{L}_{1}+\delta A,\dots,\widetilde{L}_{d}+\delta A$ on $\widetilde{X}$ as in Theorem 1.1. Using Theorem 1.1 we can now compute
(7.2) $(\widetilde{L}_{1}\cdot\ldots\cdot\widetilde{L}_{d})$ $\leq((\widetilde{L}_{1}+\delta A)\cdot\ldots\cdot(\widetilde{L}_{d}+\delta A))$
$=d!V(\Delta_{\widetilde{\nu}}(\widetilde{L}_{1}+\delta A),\dots,\Delta_{\widetilde{\nu}}(\widetilde{L}_{d}+\delta A))$
$=d!V(\Delta_{\widetilde{\nu}}(\pi^{*}L_{1}-E_{1}+\delta A),\dots,\Delta_{\widetilde{\nu}}(\pi^{*}L_{d}-E_{d}+\delta A))$
$\leq d!V(\Delta_{\widetilde{\nu}}(\pi^{*}L_{1}+\delta A),\dots,\Delta_{\widetilde{\nu}}(\pi^{*}L_{d}+\delta A)),$
where the last inequality follows from the inclusion
$\Delta_{\widetilde{\nu}}(\pi^{*}L_{i}-E_{i}+\delta A)+\widetilde{\nu}(e_{i})\subseteq\Delta_{\widetilde{\nu}}(\pi^{*}L_{i}+\delta A)$
for any $e_{i}\in H^{0}(X,E_{i})\setminus\{0\}$ which is induced in the same way as the inclusion (2.3). Note that the mixed volume is monotonically increasing and invariant under translation in each variable.
Using the continuity of Newton–Okounkov bodies and the monotonicity and multi-linearity of the mixed volume, we can choose $\delta>0$ such that
(7.3) $d!V(\Delta_{\widetilde{\nu}}(\pi^{*}L_{1}+\delta A),\dots,\Delta_{\widetilde{\nu}}(\pi^{*}L_{d}+\delta A))\leq d!V(\Delta_{\widetilde{\nu}}(\pi^{*}L_{1}),\dots,\Delta_{\widetilde{\nu}}(\pi^{*}L_{d}))+\epsilon.$
We write $\nu\colon K(X)^{\times}\to\mathbb{Z}^{d}$ for the valuation obtained from $\widetilde{\nu}$ and the isomorphism $K(X)\cong K(\widetilde{X})$ induced by $\pi$. We apply Equation (2.8) to get
(7.4) $d!V(\Delta_{\widetilde{\nu}}(\pi^{*}L_{1}),\dots,\Delta_{\widetilde{\nu}}(\pi^{*}L_{d}))=d!V(\Delta_{\nu}(L_{1}),\dots,\Delta_{\nu}(L_{d})).$
If we now combine the inequalities (7.2) and (7.3) and Equation (7.4), we obtain the inequality in (7.1) as desired.
8. Proof of Corollary 1.4
We prove Corollary 1.4 in this section. By continuity we may assume that $L,L_{1},\dots,L_{d}$ are ample $\mathbb{Q}$-line bundles. It has been shown by Saroglou–Soprunov–Zvavitch [15, Equation 6.1] that
@@
for all convex bodies $K,K_{1},\dots,K_{d}$ in $\mathbb{R}^{d}$. If $Y_{\bullet}$ is an admissible flag on $X$ satisfying (i)–(iii) of Theorem 1.2, we get by Theorem 1.2 for $k=d-1$ and by Equation (2.1)
@@
×$\quad\times V(\Delta_{Y_{\bullet}}(L_{1}),\dots,\Delta_{Y_{\bullet}}(L_{d-1}),\Delta_{Y_{\bullet}}(L))$
$\leq d(L^{d-1}\cdot L_{d})\cdot(L_{1}\cdot\ldots\cdot L_{d-1}\cdot L),$
where we used for the last inequality Theorem 1.2 and the inequality
@@
which we proved for any admissible flag $Y_{\bullet}$ and any ample line bundles $M_{1},M_{2}$ on $X$ in [17, Lemma 6.1].

### v3

@@ 4. Construction of the flag
In this section, we show the existence of an admissible flag $Y_{\bullet}$ satisfying the properties (i)-(iii) in Theorem 1.2. For simplicity, we set
@@
We proceed as in [11, Section 4]. We show by induction that for every $0\leq j\leq d-1$ there is a flag of irreducible subvarieties $X=Y_{0}\supsetneq Y_{1}\supsetneq\dots\supsetneq Y_{j}$ such that for all $0\leq i\leq j-1$ the following hold
  • $Y_{i+1}$ is an irreducible Cartier divisor on $Y_{i}$,
  • $Y_{i+1}\cap Y_{i,\mathrm{reg}}\subseteq Y_{i+1,\mathrm{reg}}$,
  • there is an $r_{i+1}\in\mathbb{Q}$ with $r_{i+1}\mathcal{O}_{Y_{i}}(Y_{i+1})\cong L^{\prime}_{i+1}|_{Y_{i}}$ as $\mathbb{Q}$-line bundles, and
  • $Y_{j}\cap X_{\mathrm{reg}}\neq\emptyset$.
This is trivial for $j=0$, and we may assume that we have constructed such a flag for some $0\leq j\leq d-2$. We must show that we can extend it by choosing $Y_{j+1}$ so that the properties hold for $i=j$ and $Y_{j+1}\cap X_{\mathrm{reg}}\neq\emptyset$.
Since $L^{\prime}_{j+1}$ is ample, $L^{\prime}_{j+1}|_{Y_{j}}$ is ample, too. Thus, we can choose a positive integer $\widetilde{r}_{j+1}\in\mathbb{Z}_{\geq 1}$ such that $\widetilde{r}_{j+1}L^{\prime}_{j+1}|_{Y_{j}}$ is very ample, and hence it induces an embedding $Y_{j}\to\mathbb{P}^{N}$. Since $\dim Y_{j}=d-j\geq 2$, we can use the following two versions of Bertini’s theorem: By [2, Theorem 1.1] any general hyperplane $H\subseteq\mathbb{P}^{N}$ intersects $Y_{j}$ in an irreducible Cartier divisor on $Y_{j}$. By [1, Corollary 2] we have $(H\cap Y_{j})_{\mathrm{reg}}\supseteq Y_{j,\mathrm{reg}}\cap H$ for any general hyperplane $H\subseteq\mathbb{P}^{N}$. Since every hyperplane $H\subseteq\mathbb{P}^{N}$ intersects $Y_{j}$, any general hyperplane $H\subseteq\mathbb{P}^{N}$ intersects the dense open subset $Y_{j}\cap X_{\mathrm{reg}}$ of $Y_{j}$. Hence, there is a hyperplane $H_{1}\subseteq\mathbb{P}^{N}$ such that $Y_{j+1}=H_{1}\cap Y_{j}$ is an irreducible Cartier divisor on $Y_{j}$, every point of $Y_{j+1}$ that is regular on $Y_{j}$ is also regular on $Y_{j+1}$, and $Y_{j+1}\cap X_{\mathrm{reg}}\neq\emptyset$. By construction we have $r_{j+1}\mathcal{O}_{Y_{j}}(Y_{j+1})\cong L^{\prime}_{j+1}|_{Y_{j}}$ for $r_{j+1}=\widetilde{r}_{j+1}^{-1}\in\mathbb{Q}$. Thus, we have proved the induction hypothesis.
To complete the flag $Y_{0}\supsetneq Y_{1}\supsetneq\dots\supsetneq Y_{d-1}$, we set $Y_{d}=\{p\}$ for any closed point $p\in Y_{d-1}\cap X_{\mathrm{reg}}$. By construction, we have $p\in Y_{j,\mathrm{reg}}$ for all $0\leq j\leq d-1$. Hence, $Y_{\bullet}$ is an admissible flag satisfying properties (i)-(iii) in Theorem 1.2 with $r_{j}=\widetilde{r}_{j}^{-1}$.
@@ 5. The Newton–Okounkov bodies in this paper
Let $L_{1},\dots,L_{d}$ be ample $\mathbb{R}$-line bundles on $X$ and $Y_{\bullet}$ an admissible flag on $X$ satisfying the corresponding properties (i)-(iii) in Theorem 1.2 with $k\geq 1$. Then it holds
@@
    As $L_{1}\equiv r_{1}\mathcal{O}_{X}(Y_{1})$, we can compute by the slice formula (2.5), Equation (2.2), and Equations (2.3) and (2.4) for $\tau\neq\mu(L_{j};Y_{1})$ that
@@
    For the first inclusion, note that we always have $\Delta_{Y_{\bullet}|Y_{1}}(L)\subseteq\Delta_{Y_{1,\bullet}}(L|_{Y_{1}})$ for any line bundle $L$ on $X$. By continuity of Newton–Okounkov bodies, the inclusion also holds for $\tau=\mu(L_{j};Y_{1})$. This completes the proof of the lemma.∎
@@
We prove Theorem 1.2 by induction on $k$. By Section 4 we only have to prove the second statement. If $k=0$, we proved in [11, Theorem 1.2] that the inclusion (2.3) is indeed an equality $\Delta_{Y_{\bullet}}(N_{1})+\Delta_{Y_{\bullet}}(N_{2})=\Delta_{Y_{\bullet}}(N_{1}+N_{2})$ if $N_{1},N_{2}\in C_{L}(M)$. Thus, the theorem follows from the multilinearity of the intersection number and the mixed volume. See also [11, Section 4]. Hence, we may assume $k\geq 1$. In particular, $d\geq 2$. Applying the induction hypothesis to $Y_{1}$ we get
$(L_{1}\cdot\ldots\cdot L_{d})$ $=r_{1}(L_{2}|_{Y_{1}}\cdot\ldots\cdot L_{d}|_{Y_{1}})$
$=r_{1}(d-1)!V(\Delta_{Y_{1,\bullet}}(L_{2}|_{Y_{1}}),\dots,\Delta_{Y_{1,\bullet}}(L_{d}|_{Y_{1}})).$
Since $L_{2},\dots,L_{d}$ are ample, we get by Equation (2.4) that
$V(\Delta_{Y_{1,\bullet}}(L_{2}|_{Y_{1}}),\dots,\Delta_{Y_{1,\bullet}}(L_{d}|_{Y_{1}}))=V(\Delta_{Y_{\bullet}|Y_{1}}(L_{2}),\dots,\Delta_{Y_{\bullet}|Y_{1}}(L_{d})).$
@@
where we consider the convex bodies in the mixed volume on the right-hand side as convex bodies in $\mathbb{R}^{d-1}\cong\{0\}\times\mathbb{R}^{d-1}$. By Lemmas 3.1 and 5.1 we get
@@
We prove Corollary 1.3 in this section. By continuity, we may assume that $L,L_{1},\dots,L_{d}$ are ample $\mathbb{Q}$-line bundles. It has been shown by Saroglou–Soprunov–Zvavitch [9, Equation 6.1] that
@@
for all convex bodies $K,K_{1},\dots,K_{d}$ in $\mathbb{R}^{d}$. If $Y_{\bullet}$ is an admissible flag on $X$ satisfying (i)-(iii) of Theorem 1.2, we get by Theorem 1.2 for k=d1k=d-1 and by Equation (2.1)
@@
$\quad\times V(\Delta_{Y_{\bullet}}(L_{1}),\dots,\Delta_{Y_{\bullet}}(L_{d-1}),\Delta_{Y_{\bullet}}(L))$
$\leq d(L^{d-1}\cdot L_{d})\cdot(L_{1}\cdot\ldots\cdot L_{d-1}\cdot L).$
For the last inequality, we used Theorem 1.2 together with the inequality
@@
which we proved for any admissible flag $Y_{\bullet}$ and any ample line bundles $M_{1},M_{2}$ on $X$ in [11, Lemma 6.1].